find the equation of the line that contains the point (3,2) and is perpendicular to the line 3x+y= -3
thanks for the help!

Answer by mgmoeab(37)

The line you're looking for passes through the point (3,2) and is perpendicular to the line 3x+y= -3
You're looking for the equation of the line:
y-Y1 = m(x-X1) since you are given point (3,2)
you have
a) y - 2 = m( x - 3)

Now, you only have to look for m (the slope)
-Since the line is perpendicular to 3x+y= -3, then the product of the slopes of two lines will –1.
-Since the slope of the line 3x+y= -3 is m = -3, then the slope of the line in question is m= 1/3
The final step is to plug in the slope we found and the rest depends if you are asked to solve for y (b), or leave the equation in point-slope-form (a)
(a) y - 2 = 1/3( x - 3)
(b)y = 1/3x + 1
HOPE IT HELPS